# Convergence Rates of Constrained Expected Improvement

### v1

@@ Appendix D Example details
$f^{*}=0.25.$
The objective, constraint functions and the optimal $f$ of example 4 is given below.
@@
$f^{*}=0.6.$
The objective, constraint functions and the optimal $f$ of example 5 is given below.
@@
The objective, constraint functions and the optimal $f$ of example 6 is given below.
@@
$f^{*}=-3.32.$
The objective, constraint functions and the optimal $f$ of example 7 is given below.
$f(\mathbf{x})=100(x_{2}-x_{1}^{2})^{2}+(1-x_{1})^{2}$ (101)
@@
The contour plots of example 1, 2, 3, 4, and 7 are given in Figure 2345, and 6.
Figure 2: Contour plots for the objective function (left) and constraint function (right) for example 1. The feasible region is marked on the plots, along with the global optimum (* sign).
Figure 3: Contour plots for the objective function (left) and constraint function (right) for example 2. The feasible region is marked on the plots, along with the global optimum (* sign).
Figure 4: Contour plots for the objective function (left) and constraint function (right) for example 3. The feasible region is marked on the plots, along with the global optimum (* sign).
Figure 5: Contour plots for the objective function (left) and the two constraint functions (middle and right) for example 4. The feasible region is marked with black line on the objective contour, along with the global optimum (* sign).
Figure 6: Contour plots for the objective function (left) and constraint function (right) for example 7. The feasible region is marked on the plots, along with the global optimum (* sign).

### v2

@@ Appendix D Test problems
[…]
The objective, constraint functions and the optimal $f$ of Problem 2 is given below.
@@
[…]
The objective, constraint functions and the optimal $f$ of Problem 3 is given below.
@@
The objective, constraint functions and the optimal $f$ of Problem 4 is given below.
@@
[…]
Figure 3: Contour plots for the objective function (left) and constraint function (right) for Problem 1. The infeasible region is marked on the plots. The global optimum is marked with a star sign.
Figure 4: Contour plots for the objective function (left) and the two constraint functions (middle and right) for Problem 2. The infeasible region is marked with black line on the objective contour. The global optimum is marked with a star sign.
Figure 5: Contour plots for the objective function (left) and constraint function (right) for Problem 5. The infeasible region is marked on the plots. The global optimum is marked with a star sign.
The objective, constraint functions and the optimal $f$ of Problem 5 is given below.
[…]
@@
The contour plots of Problem 1, 2, and 5 are given in Figure 34, and 5.